The length of a rectangular garden is two meters longer than twice its width. The perimeter is 46 meters. Find the dimensions of the garden.
Answer by addingup(3677)

2L+2W= 46 and L= 2W+2
Substitute for L:
2(2W+2)+2W= 46
4W+4+2W= 46
6W+4= 46
6W = 42
W = 7
L = 2(7)+2 = 16
The garden is 7 wide by 16 long
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2(7)+2(16)= 46
14+32 = 46 We have the right answer
J